What is the impedance (as a function of frequency) of a 25nF capacitor in series with a 100 Ω resistor?

In series?
Answer #1

Ans. The capacitor and the resistor are in series, so the impedance Z = R + Xc

Xc = capacitive reactance and R = 100 Ohms

2π fc C = 25 nF = 25 * 10-9 F So. X.- 1 637 * 106 The impedance Z = 100+ б.3( * 10.2 2Tf 25 10-5f